# Planning in Branch-and-Bound: Model-Based Reinforcement Learning for Exact Combinatorial Optimization

### v1

@@ I   Further computational results
To contextualize the results presented in Section 5, we provide additional discussion on the impact of leveraging the PlanB&B model to improve branching decisions. In our experiments, we observe that outside of the MIS benchmark, the value head $\bar{\mathbf{v}}$ of the prediction network $f$ struggles to generalize to higher-dimensional MILPs. We attribute this to out-of-distribution effects: the value network is trained on instances with structurally smaller subtrees than those encountered in the transfer benchmarks. Similar limitations are observed in prior RL baselines: PG-tMDP demonstrates stronger generalization than both DQN agents, likely due to its reliance on a policy network rather than a $Q$-network to make branching decisions. Fortunately, this limitation does not prevent the policy network $\mathbf{p}$ from producing efficient branching strategies on transfer instances, as shown in Table 1. However, it does hinder the ability of PlanB&B to refine branching decisions at inference time using its learned model, as the inaccuracies of the value predictions undermine planning. This limitation could potentially be addressed in future work by gradually increasing the dimensionality of the instances solved during training by the MILP actors, thereby allowing the value network to progressively adapt to larger B&B trees such as the one encountered in transfer benchmarks.

### v2

@@ I   Further computational results
To contextualize the results presented in Section 5, we provide additional discussion on the impact of leveraging the PlanB&B model to improve branching decisions. In our experiments, we observe that outside of the MIS benchmark, the value head $\bar{\mathbf{v}}$ of the prediction network $f$ struggles to generalize to higher-dimensional MILPs. We attribute this to out-of-distribution effects: the value network is trained on instances with structurally smaller subtrees than those encountered in the transfer benchmarks. Similar limitations are observed in prior RL baselines: PG-tMDP demonstrates stronger generalization than both DQN agents, likely due to its reliance on a policy network rather than a $Q$-network to make branching decisions. Fortunately, this limitation does not prevent the policy network $\mathbf{p}$ from producing efficient branching strategies on transfer instances, as shown in Table 1. However, it does hinder the ability of PlanB&B to refine branching decisions at inference time using its learned model, as the inaccuracies of the value predictions undermine planning. This is illustrated in Figure 7 on combinatorial auction instances: improved node performance can be achieved by leveraging model-based planning on test instances, but not on transfer instances. This limitation could potentially be addressed in future work by gradually increasing the dimensionality of the instances solved during training by the MILP actors, thereby allowing the value network to progressively adapt to larger B&B trees such as the one encountered in transfer benchmarks.
(a) Test instances
(b) Transfer instances
Figure 7: Policy improvement associated with increased simulation budget over the CA benchmark.